OpenStudy (anonymous):

what is the value of the expression 4x (3x under the radicand) + 5x (8x under the radicand)?

OpenStudy (anonymous):

4x\[\sqrt{3x}\]+5x\[\sqrt{?8x}\]

OpenStudy (anonymous):

4xsqrt{3x}+5x\sqrt{8x}\]

OpenStudy (anonymous):

sorry, i dont understand how to type the symbols